OpenStudy (anonymous):

find the circumference of the circle whose area is 16times the area of the cirlce with diameter 1.4cm

OpenStudy (anonymous):

@Jhannybean

OpenStudy (anonymous):

@mathmath333

OpenStudy (mathmath333):

\(\large\tt \begin{align} \color{black}{ A_c=16A_1 }\end{align}\) u agree

OpenStudy (anonymous):

yes

OpenStudy (mathmath333):

\(\large\tt \begin{align} \color{black}{ A_c=16A_1\\~\\ \pi (r_c)^2=16\pi (r_1)^2\\~\\ }\end{align}\)

OpenStudy (mathmath333):

is that ok

OpenStudy (anonymous):

yes

OpenStudy (mathmath333):

\(\large\tt \begin{align} \color{black}{ (r_c)^2=16( r_1)^2\\~\\ r_c=4r_1 }\end{align}\) ok

OpenStudy (anonymous):

yes

OpenStudy (mathmath333):

\(\large\tt \begin{align} \color{black}{ r_c=4r_1\\~\\ r_c=2(2r_1)\\~\\ r_c=2(d_1)\\~\\ r_c=2(1.4)\\~\\ }\end{align}\) is thatok

OpenStudy (anonymous):

yes

OpenStudy (mathmath333):

\(\large\tt \begin{align} \color{black}{ r_c=2(1.4)\\~\\ r_c=2.8\\~\\ 2\pi r_c=2\pi\times 2.8\\~\\ C=2\pi \times 2.8\\~\\ }\end{align}\)

OpenStudy (anonymous):

is it 22 x 0.8?

OpenStudy (mathmath333):

\(C=2\times \dfrac{22}{7}\times 2.8\\~\\\)

OpenStudy (anonymous):

i did that i got 44 x 0.8??